OpenStudy (anonymous):

square 3y+4=4 solution for y =?

OpenStudy (amistre64):

can you retype the question, its kinda confusing the way it is

OpenStudy (anonymous):

\[\sqrt{3y+4=4} solve for y\]

OpenStudy (anonymous):

Are you sure the square root sign doesn't stop at the = sign?

OpenStudy (anonymous):

To get rid of a square root you square both sides. So square both sides and what do you get?

OpenStudy (anonymous):

I have no clue

OpenStudy (amistre64):

sqrt(3y+4) = 4 ?? perhaps

OpenStudy (amistre64):

3y+4 = 16 3y = 12 y = 4 then

OpenStudy (anonymous):

oh ok that makes since

OpenStudy (anonymous):

so my problem is sqrt (3y+6)=6 so the answer would be complete like this 3y+6= 24 3y=18 y=6

OpenStudy (anonymous):

b^{3}=-64 I completed by dividing -64 by 3 I got the answer -21.33333 Am I completely wrong about this problem

OpenStudy (anonymous):

\[\sqrt{h/2}\] 3 feet from the ground determine the seconds the person in the air